A tank is filled by a pipe in 18 hours and emptied by another pipe in 12 hours. How long will it take to empty the full tank if both the pipes are kept open simultaneously?
Answer by ptaylor(2198)

Let x=amount of time it takes to empty the full tank when both pipes are open
The fill pipe fills at the rate of 1/18 tank per hour
The drain pipe drains at the rate of 1/12 tank per hour
When both pipes are kept open, they drain at the rate of 1/12-1/18=3/36-2/36=1/36 tank per hour. Sooo, our equation to solve is:
(1/36)*x=1 (tank, that is)
x=36 hrs----time to empty the full tank when both pipes are open.
Hope this helps---ptaylor